OpenStudy (anonymous):

I need help with multiplying radicals, can someone pull up with an example problem and help me?

OpenStudy (tkhunny):

\(\sqrt{x}\cdot\sqrt{x} = x\) \(\sqrt{x^{2}} = |x|\) I'm just getting started. Maybe a little more clue would help.